Hunter invested $750 in an account paying an interest rate of 6 5/8% compounded continuously. London invested $750 in an account paying an interest rate of 6 1/2% compounded daily. After 18 years, how much more money would Hunter have in his account then London, to the nearest dollar?
Answer by Boreal(15235)

a. P=750*e^(rt)=750*e^(.06625*18)
=750*e^1.1925
=$2471.48 or $2471
-
b. P=750(1+(0.0650/365))^(365*18)
=$2416.24 or $2416
-
The difference is $55.24 or $55.